OpenStudy (anonymous):

Identify the correct slope and y-intercept of the equation 5x - 3y = -6

OpenStudy (anonymous):

slope=5

OpenStudy (anonymous):

what is the y=mx+b form

OpenStudy (anonymous):

y inye.=5(0)-3y=-6 -3y=-6 -6/-3 y=3 that is the slope intersect form..

OpenStudy (anonymous):

x=0

OpenStudy (anonymous):

what are the points

OpenStudy (anonymous):

@mulla slope intercept form is y=5x/3 +2 ..so slope is 5/3 and y intercept 2

OpenStudy (anonymous):

no 3y=5x+6 right?

OpenStudy (anonymous):

y=mx+c that is the slope intercept form

OpenStudy (anonymous):

yaa but its 3y in this question right???????

OpenStudy (anonymous):

the answer choices are 1. slope = -5; y-intercept = (0, -6) 2. slope=5/3;y intercept=(0,2) 3. slope= -3/5 y intercept=(0,2) 4. slope=5/3 y intercept=(0,-6)

OpenStudy (anonymous):

ya so divide by 3 to make it in standard form..which is y=5x/3 +2 ..so slope is 5/3 and y intercept 2or (0,2)

OpenStudy (anonymous):

i am going with option 1